OpenStudy (anonymous):

Given F(x) shown below, complete the equation for the inverse of F(x). If necessary, use the slash mark (/) for the division symbol. F(x)=5x+4 f^-1(y)=_______

hero (hero):

Are you sure that's what the question is?

OpenStudy (anonymous):

would you -4 to both sides and get y/1?

hero (hero):

It should be asking what's \(F^{-1}(x)\)

OpenStudy (anonymous):

and yes im sure thats what the question says

OpenStudy (anonymous):

yes you are correct hero

hero (hero):

I know I am

hero (hero):

To get the inverse, let y = \(F(x)\)

hero (hero):

y = 5x + 4

hero (hero):

Then swap x and y: x = 5y + 4

hero (hero):

Then isolate y

hero (hero):

x - 4 = 5y \[\frac{x - 4}{5} = y\] Then replace y with \(F^{-1}(x)\)

hero (hero):

\[\frac{x - 4}{5} = F^{-1}(x)\] I know that for you the steps might be confusing, but they are always the same.

OpenStudy (anonymous):

damn im confused

hero (hero):

There's nothing confusing about it.

hero (hero):

Where are you confused?

OpenStudy (anonymous):

on the work do i continue to find the answe r

OpenStudy (anonymous):

hero still theer

hero (hero):

Yes, I'm here. I showed you all the steps.